# GraphPerf-RT: A Graph-Driven Performance Model for Hardware-Aware Scheduling of OpenMP Codes

### v1

@@ 1 Introduction
Heterogeneous embedded Systems-on-Chip (SoCs) combine high-performance and energy-efficient cores with Dynamic Voltage and Frequency Scaling (DVFS). This architecture has created unprecedented opportunities for deploying parallel applications in resource-constrained environments such as autonomous vehicles, robotics, and edge computing. OpenMP, the dominant shared-memory parallel programming model, enables developers to express task-level parallelism through pragma-based annotations that the runtime system maps to available hardware resources. However, achieving optimal performance on these heterogeneous platforms requires careful scheduling decisions that balance execution time, energy efficiency, and thermal constraints. This presents a challenging optimization problem that current approaches fail to address comprehensively. Performance prediction for parallel workloads has traditionally relied on analytical models, simulation-based methods, and machine learning surrogates. Analytical models grounded in queueing theory and scheduling bounds such as Brent’s theorem provide theoretical insights but require simplifying assumptions that break down under irregular control flow and cache contention (graham1969bounds; brent1974parallel). Simulation-based approaches offer high fidelity but incur prohibitive runtime overhead for online scheduling decisions. Machine learning surrogates, including Graph Neural Networks (GNNs), can learn complex input-output mappings from data (velickovic2017graph). However, existing approaches such as ProGraML (cummins2021programl) and Ithemal (mendis2019ithemal) operate on static intermediate representations without incorporating runtime context, thermal state, or DVFS configurations that critically affect embedded system performance.
@@ 2 Related Work
Performance modeling for parallel systems spans analytical queueing/Petri-net models and simulators (eager1983bound; marsan1984class), then statistical and neural surrogates (lee2007accurate; ipek2006efficiently; wang2018machine; chen2016xgboost). Black-box predictors (counters, instruction mixes) improve accuracy but lose structural signals and transfer. Learned models like Ithemal (mendis2019ithemal) use hierarchical LSTMs to predict x86 basic-block throughput from instruction sequences, achieving high accuracy on microarchitectural latency but operating on static assembly without capturing runtime parallelism, inter-task dependencies, or device/thermal state. Runtime autotuners (Apollo) adapt to input-dependent performance variations (ramadan2024), yet remain kernel-focused. Our focus differs: we model OpenMP task DAGs on embedded SoCs, explicitly encoding dependencies, per-core DVFS, and thermal context to capture cross-layer effects needed for real-time decisions.
@@
Graph neural networks enable learning on structured program and system graphs (wu2020comprehensive; velickovic2017graph). ProGraML (cummins2021programl) constructs a unified graph representation from LLVM IR by combining control-flow, data-flow, and call edges, achieving strong results on compiler optimization tasks such as device mapping and thread coarsening; critically, it operates exclusively on static intermediate representation without runtime context, hardware state, or thermal conditions. Similarly, Ithemal (mendis2019ithemal) predicts microarchitectural latency from static instruction sequences but lacks parallelism structure, inter-task dependencies, or device state. Heterogeneous Graph Transformer (HGT) (hu2020heterogeneous) extends graph attention to multiple node and edge types, providing a natural baseline for our heterogeneous task-resource graphs. However, HGT and similar architectures lack three critical capabilities for our embedded scheduling setting: (1) explicit edge decomposition that separates structural dependencies from execution features, (2) integration of thermal and DVFS context as part of the graph representation, and (3) evidential regression heads that produce calibrated uncertainty estimates for risk-aware decision-making. GNNs have also predicted hardware-dependent performance for ML models (shi2022graph) and optimization effects (brauckmann2020compiler). Compiler cost models (Ansor, MetaSchedule, ROLLER) learn evaluators for tensor kernels and improve cross-device transfer (zheng2020ansor; shao2022metaschedule; zhu2022roller; zhai2023tlp; zhai2024tlm), but operate below application-level task graphs and rarely encode runtime DVFS/thermal state. Recent work on transferable hardware embeddings includes Silhouette (papon2022silhouette), which learns performance-conscious CPU embeddings for cross-platform prediction, while Glimpse (ahn2022glimpse) develops mathematical encodings of hardware specifications for neural compilation, provides an alternative to hand-crafted device sheets and could potentially enhance resource node features in our framework. However, these approaches focus on static kernel-level prediction without integrating dynamic per-core DVFS, thermal headroom, or application-level task DAG structure that we require for OpenMP scheduling under thermal constraints. In contrast, we construct a heterogeneous graph with task and resource nodes, typed edges with execution features, and CFG-informed encoders to improve zero-shot and few-shot generalization across benchmarks and boards while explicitly fusing runtime device/thermal context into the representation.
@@
Hardware-aware ML includes NAS with device constraints and cross-platform predictors (cai2018proxylessnas), and performance modeling across CPU/heterogeneous configurations (walker2018accurate). Energy/thermal control spans heuristic/optimization DVFS and RL (xie2021survey). Utilization-driven governors can mis-scale frequency by conflating compute and stalls (hebbar2022pmu; lin2023workload). Model-free RL adapts but is sample- and heat-intensive on small boards (kim2021ztt). Model-based control reduces sampling via learned dynamics and planning (moerland2023model). Our approach is complementary: a calibrated, graph-grounded evaluator supplies fast, hardware-aware rollouts to search/MARL under DVFS and thermal constraints.
@@
Uncertainty is critical for safe scheduling. Bayesian and ensemble methods provide confidence but add inference cost (lakshminarayanan2017simple; kendall2017uncertainties). Evidential learning yields calibrated intervals without sampling (sensoy2018evidential; amini2020deep). We adopt evidential regression heads to expose confidence on makespan, energy, and counters and to gate synthetic rollouts, aligning with on-board, real-time planning.
@@ 4.2.2 Heterogeneous GAT Layers
where $\mathcal{R}=\{TT,TR,RR,RM\}$ is the set of edge types, $\mathcal{N}_{r}(v)$ is the neighborhood of $v$ under edge type $r$, $\alpha_{uv,r}^{(l)}$ is the attention coefficient, $\mathbf{W}_{r}^{(l)}$ is a type-specific linear transformation matrix, and $\sigma$ is a non-linearity (e.g., ELU (clevert2015elu)). Residual connections are added for deeper models to mitigate vanishing gradients.
@@ 4.4 Training Procedure
Stage 3: Calibration. Finally, the model undergoes calibration using temperature scaling (guo2017calibration) or isotonic regression on held-out data to ensure that uncertainty estimates are well-calibrated across different performance metrics and operating conditions (e.g., expected calibration error <0.05<0.05). This involves optimizing a post-hoc scalar per target to align predicted confidences with empirical accuracies.
@@ 5.2 Baselines and Metrics
Heterogeneous Graph Transformer (HGT) (hu2020heterogeneous) represents a state-of-the-art heterogeneous GNN that uses type-aware attention over our node and edge schema as defined in Section 3.2. This baseline isolates the contribution of our evidential heads and runtime context integration, as HGT lacks evidential uncertainty quantification, thermal and utilization context in node features, and our CFG-derived task features.
@@ H.1 DAG Structural Metrics
Critical Path Length (Span). For a task DAG $G=(V,E)$ with vertex weights $w:V\rightarrow\mathbb{R}^{+}$ representing task execution times, the span $T_{\infty}$ is the length of the longest weighted directed path from any source to any sink (graham1969bounds; cormen2009introduction):
@@
Under greedy scheduling with $P$ processors, Brent’s theorem (brent1974parallel) bounds makespan as $T_{P}\leq T_{1}/P+T_{\infty}$. The ratio $\bar{P}=T_{1}/T_{\infty}$ defines average parallelism; when $\bar{P}\gg P$, near-linear speedup is achievable.
@@
While standard learning theory provides generalization bounds via Rademacher complexity, these bounds scale poorly with network depth and width for GNNs (scarselli2009graph). Message-passing GNNs achieve universal approximation under certain conditions, though expressiveness is bounded by the Weisfeiler-Leman graph isomorphism test (xu2019powerful; morris2019weisfeiler). We validate generalization empirically in Section 5.
@@ H.2 Graph Complexity Measures for DAGs
DAG Width. The width at level $\ell$ is the number of tasks at topological distance $\ell$ from source nodes (i.e., tasks $v$ where all paths from sources to $v$ have length $\ell$). Maximum width bounds parallelism under ideal scheduling (blumofe1999scheduling; kwok1999static). Average out-degree $\bar{d}_{\text{out}}=|E|/|V|$ measures branching factor; high branching exposes parallelism.
@@ H.3 Evidential Uncertainty Theory
Normal-Inverse-Gamma (NIG) Prior. Evidential regression (amini2020deep) places a higher-order prior over Gaussian likelihood parameters. For target $y\in\mathbb{R}$:
@@
Uncertainty Decomposition. Following amini2020deep, the total predictive variance decomposes into aleatoric and epistemic components:
@@
Evidential Loss. The model outputs $(\gamma,\nu,\alpha,\beta)$ in a single forward pass. Training minimizes the negative log-marginal likelihood with evidence regularization (amini2020deep, Eq. 10–11):
@@
Computational Advantage. Bayesian approaches (MC-Dropout, deep ensembles) achieve similar calibration but require 10–100 forward passes (gal2016dropout; lakshminarayanan2017simple). Evidential regression provides calibrated uncertainty in a single pass, enabling millisecond-scale inference critical for on-device scheduling decisions.

### v2

@@ 1 Introduction
Heterogeneous embedded Systems-on-Chip (SoCs) combine high-performance and energy-efficient cores with Dynamic Voltage and Frequency Scaling (DVFS). This architecture has created unprecedented opportunities for deploying parallel applications in resource-constrained environments such as autonomous vehicles, robotics, and edge computing. OpenMP, the dominant shared-memory parallel programming model, enables developers to express task-level parallelism through pragma-based annotations that the runtime system maps to available hardware resources. However, achieving optimal performance on these heterogeneous platforms requires careful scheduling decisions that balance execution time, energy efficiency, and thermal constraints. This presents a challenging optimization problem that current approaches fail to address comprehensively. Performance prediction for parallel workloads has traditionally relied on analytical models, simulation-based methods, and machine learning surrogates. Analytical models grounded in queueing theory and scheduling bounds such as Brent’s theorem provide theoretical insights but require simplifying assumptions that break down under irregular control flow and cache contention [13, 5]. Simulation-based approaches offer high fidelity but incur prohibitive runtime overhead for online scheduling decisions. Machine learning surrogates, including Graph Neural Networks (GNNs), can learn complex input-output mappings from data [39]. However, existing approaches such as ProGraML [10] and Ithemal [26] operate on static intermediate representations without incorporating runtime context, thermal state, or DVFS configurations that critically affect embedded system performance. Complementary approaches address related challenges through hierarchical multi-agent DVFS scheduling [31], zero-shot LLM-guided allocation [30], statistical feature-aware task allocation [32], and flow-augmented few-shot RL [33].
@@ 2 Related Work
Performance modeling for parallel systems spans analytical queueing/Petri-net models and simulators [11, 25], then statistical and neural surrogates [23, 17, 41, 7]. Black-box predictors (counters, instruction mixes) improve accuracy but lose structural signals and transfer. Learned models like Ithemal [26] use hierarchical LSTMs to predict x86 basic-block throughput from instruction sequences, achieving high accuracy on microarchitectural latency but operating on static assembly without capturing runtime parallelism, inter-task dependencies, or device/thermal state. Runtime autotuners (Apollo) adapt to input-dependent performance variations [34], yet remain kernel-focused. Our focus differs: we model OpenMP task DAGs on embedded SoCs, explicitly encoding dependencies, per-core DVFS, and thermal context to capture cross-layer effects needed for real-time decisions.
@@
Graph neural networks enable learning on structured program and system graphs [42, 39]. ProGraML [10] constructs a unified graph representation from LLVM IR by combining control-flow, data-flow, and call edges, achieving strong results on compiler optimization tasks such as device mapping and thread coarsening; critically, it operates exclusively on static intermediate representation without runtime context, hardware state, or thermal conditions. Similarly, Ithemal [26] predicts microarchitectural latency from static instruction sequences but lacks parallelism structure, inter-task dependencies, or device state. Heterogeneous Graph Transformer (HGT) [16] extends graph attention to multiple node and edge types, providing a natural baseline for our heterogeneous task-resource graphs. However, HGT and similar architectures lack three critical capabilities for our embedded scheduling setting: (1) explicit edge decomposition that separates structural dependencies from execution features, (2) integration of thermal and DVFS context as part of the graph representation, and (3) evidential regression heads that produce calibrated uncertainty estimates for risk-aware decision-making. GNNs have also predicted hardware-dependent performance for ML models [38] and optimization effects [4]. Compiler cost models (Ansor, MetaSchedule, ROLLER) learn evaluators for tensor kernels and improve cross-device transfer [47, 37, 48, 45, 46], but operate below application-level task graphs and rarely encode runtime DVFS/thermal state. Recent work on transferable hardware embeddings includes Silhouette [29], which learns performance-conscious CPU embeddings for cross-platform prediction, while Glimpse [1] develops mathematical encodings of hardware specifications for neural compilation, provides an alternative to hand-crafted device sheets and could potentially enhance resource node features in our framework. However, these approaches focus on static kernel-level prediction without integrating dynamic per-core DVFS, thermal headroom, or application-level task DAG structure that we require for OpenMP scheduling under thermal constraints. In contrast, we construct a heterogeneous graph with task and resource nodes, typed edges with execution features, and CFG-informed encoders to improve zero-shot and few-shot generalization across benchmarks and boards while explicitly fusing runtime device/thermal context into the representation.
@@
Hardware-aware ML includes NAS with device constraints and cross-platform predictors [6], and performance modeling across CPU/heterogeneous configurations [40]. Energy/thermal control spans heuristic/optimization DVFS and RL [43]. Utilization-driven governors can mis-scale frequency by conflating compute and stalls [15, 24]. Model-free RL adapts but is sample- and heat-intensive on small boards [19]. Model-based control reduces sampling via learned dynamics and planning [27]. Recent approaches address these gaps through hierarchical multi-agent DVFS scheduling [31], statistical feature-aware task allocation [32], zero-shot LLM-guided allocation [30], and flow-augmented few-shot RL [33]. Our approach is complementary: a calibrated, graph-grounded evaluator supplies fast, hardware-aware rollouts to search/MARL under DVFS and thermal constraints.
@@
Uncertainty is critical for safe scheduling. Bayesian and ensemble methods provide confidence but add inference cost [22, 18]. Evidential learning yields calibrated intervals without sampling [36, 2]. We adopt evidential regression heads to expose confidence on makespan, energy, and counters and to gate synthetic rollouts, aligning with on-board, real-time planning.
@@ 4.2.2 Heterogeneous GAT Layers
where $\mathcal{R}=\{TT,TR,RR,RM\}$ is the set of edge types, $\mathcal{N}_{r}(v)$ is the neighborhood of $v$ under edge type $r$, $\alpha_{uv,r}^{(l)}$ is the attention coefficient, $\mathbf{W}_{r}^{(l)}$ is a type-specific linear transformation matrix, and $\sigma$ is a non-linearity (e.g., ELU [8]). Residual connections are added for deeper models to mitigate vanishing gradients.
@@ 4.4 Training Procedure
Stage 3: Calibration. Finally, the model undergoes calibration using temperature scaling [14] or isotonic regression on held-out data to ensure that uncertainty estimates are well-calibrated across different performance metrics and operating conditions (e.g., expected calibration error <0.05<0.05). This involves optimizing a post-hoc scalar per target to align predicted confidences with empirical accuracies.
@@ 5.2 Baselines and Metrics
Heterogeneous Graph Transformer (HGT) [16] represents a state-of-the-art heterogeneous GNN that uses type-aware attention over our node and edge schema as defined in Section 3.2. This baseline isolates the contribution of our evidential heads and runtime context integration, as HGT lacks evidential uncertainty quantification, thermal and utilization context in node features, and our CFG-derived task features.
@@ H.1 DAG Structural Metrics
Critical Path Length (Span). For a task DAG $G=(V,E)$ with vertex weights $w:V\rightarrow\mathbb{R}^{+}$ representing task execution times, the span $T_{\infty}$ is the length of the longest weighted directed path from any source to any sink [13, 9]:
@@
Under greedy scheduling with $P$ processors, Brent’s theorem [5] bounds makespan as $T_{P}\leq T_{1}/P+T_{\infty}$. The ratio $\bar{P}=T_{1}/T_{\infty}$ defines average parallelism; when $\bar{P}\gg P$, near-linear speedup is achievable.
@@
While standard learning theory provides generalization bounds via Rademacher complexity, these bounds scale poorly with network depth and width for GNNs [35]. Message-passing GNNs achieve universal approximation under certain conditions, though expressiveness is bounded by the Weisfeiler-Leman graph isomorphism test [44, 28]. We validate generalization empirically in Section 5.
@@ H.2 Graph Complexity Measures for DAGs
DAG Width. The width at level $\ell$ is the number of tasks at topological distance $\ell$ from source nodes (i.e., tasks $v$ where all paths from sources to $v$ have length $\ell$). Maximum width bounds parallelism under ideal scheduling [3, 21]. Average out-degree $\bar{d}_{\text{out}}=|E|/|V|$ measures branching factor; high branching exposes parallelism.
@@ H.3 Evidential Uncertainty Theory
Normal-Inverse-Gamma (NIG) Prior. Evidential regression [2] places a higher-order prior over Gaussian likelihood parameters. For target $y\in\mathbb{R}$:
@@
Uncertainty Decomposition. Following Amini et al. [2, Eq. 9], the total predictive variance decomposes into aleatoric and epistemic components:
@@
Evidential Loss. The model outputs $(\gamma,\nu,\alpha,\beta)$ in a single forward pass. Training minimizes the negative log-marginal likelihood with evidence regularization [2, Eq. 10–11]:
@@
Computational Advantage. Bayesian approaches (MC-Dropout, deep ensembles) achieve similar calibration but require 10–100 forward passes [12, 22]. Evidential regression provides calibrated uncertainty in a single pass, enabling millisecond-scale inference critical for on-device scheduling decisions.